The Justice League owns shares in Wayne Manor, in where they own 55% of Wayne Manor. This means that they are not able to consolidate and combine both of their financial statements. The Justice League received annual dividends from Wayne Manor in the amount of $1,500,000. If The Justice League’s marginal tax rate is 35% and if the exclusion on inter-company dividends is 70%:

What is the effective tax rate on the inter-company dividends, and how much will the net dividends be, that The Justice League will receive?

10.5%; $1,342,500

10.5%; $1,185,000

10.2%; $1,715,500

10.2%; $1,472,000

10.2%; $1,250,500
Answer #1

Tax will be levied on 1,500,000*(100-70)% =$450,000

Tax = 450,000*35% =$157,500

Effective tax rate = 35%*30% =10.5%

Net dividend = 1,500,000 - 157,500

=$1,342,500

Hence, the answer is

10.5%, $1,342,500